An employee works 39 regular work week in August 2018. He was hired six years ago, earns a salary of $116.500/year and is exempt from the overtime provisions of the FLSA. to date he has no compensatin beyound his annual salary. He ask the employer to withhold 12% of gross pay, which is to be contributed to a 403b plan. Taxable income for federal income tax withholding? Taxable income for Social Security tax? Taxable income for medicare tax?
Answer #1

Salary = $ 116,500/year. 12% deduction for 403 (b) Plan

1. Taxable income for federal income tax withholding = 116,500 - 12% of 116,500 = 116,500 - 13,980 = $ 102,520

$ 102,520 is the annual amount. Divide this by 52 to get weekly amount i.e. 102,520/52 = $ 1,971.54

2. Taxable income for Social Security Tax = $ 116,500 or 116,500/52 = $ 2,240.38

3. Taxable income for medicare tax = $ 116,500 or 116,500/52 = $ 2,240.38